An aqueous solution contains 0.87 g/L of an organic compound. The osmotic pressure of this solution is 83 mmHg at 29°C. What is the molar mass of the solute?
1
Expert's answer
2020-04-22T09:17:23-0400

Posm=CM*R*T; [kPa]

Posm=83 mmHg =11,1 kPa

R=8.314 J/(mol*K)

T=29°C=273+29=302 K

CM=n/V=m/(Mr*V)


If V=1, m=0.87 g, then

CM=0.87/Mr


Posm=0.87/(Mr)*R*T;

Mr=0.87*8.314*302/11.1=196.8=197 g/mol

Answer: the molar mass of the solute is 197 g/mol